A sample of sulfur dioxide occupies a volume of 549 mL at 47° C and 779 mm Hg. What volume will the sulfur dioxide occupy in ml at 2° C and 0.8 atm? Round to the nearest two decimal places.


1
Expert's answer
2021-02-23T04:27:34-0500

"V_1 = 549 \;mL \\

T_1 = 47 + 273 = 320 \;K \\

P_1 = 779 \;mmHg = 1.02 \;atm\\

T_2 = 2 + 273 = 275 \;K \\

P_2 = 0.8 \;atm \\

\frac{P_1 \times V_1}{T_1} = \frac{P_2 \times V_2}{T_2} \\

V_2 = \frac{P_1 \times V_1 \times T_2}{T_1 \times P_2} \\

= \frac{1.02 \times 549 \times 275}{320 \times 0.8} \\

= 601.54 \;mL"

Answer: 601.54 mL